### v1

# Towards an exact electronic quantum many-body treatment of Kondo correlation in magnetic impurities

Abstract
The Kondo effect is a prototypical quantum phenomenon arising from the interaction between localized electrons in a magnetic impurity and itinerant electrons in a metallic host. Although it has served as the testing ground for quantum many-body methods for decades, the precise description of Kondo physics with material specificity remains challenging. Here, we present a systematic ab initio approach to converge towards an exact zero-temperature electronic treatment of Kondo correlations. Across a series of $3d$ transition metals, we extract Kondo temperatures matching the subtle experimental trends, with an accuracy far exceeding that of standard models. We further obtain microscopic insight into the origin of these trends. More broadly, we demonstrate the possibility to start from fully ab initio many-body simulations and push towards the realm of converged predictions.
Introduction
The Kondo system of a magnetic impurity embedded in a non-magnetic metallic host is a foundational quantum many-body materials problem [Hewson1993a, Kondo1964, Knorr2002, Nevidomskyy2009, Pruser2011, Kouwenhoven2001, lucignano2009kondo]. As the temperature is decreased below a characteristic Kondo temperature $T_{\mathrm{K}}$, the impurity moment is screened by the conduction electrons, forming a many-electron singlet state that manifests as a sharp resonance in the local density of states. Although the physics involves many-electron correlations in a bulk material, a qualitative understanding was achieved decades ago through the Anderson impurity model (AIM) [Anderson1961] and via Wilson’s numerical renormalization group (NRG) solution of the Kondo spin problem [Wilson1975b]. However, while deeply insightful, these model frameworks are limited in their quantitative precision due to the uncertainties in the specific model parameters and the neglect of higher-energy electronic degrees of freedom. In this work, we show that one can now describe Kondo physics without ever simplifying to such models, starting only with the ab initio many-electron Hamiltonian of the full material. In so doing, we present a precise material-specific treatment of the phenomenon, which can in principle be converged towards a numerically exact solution.
The challenge of the Kondo problem stems from the simultaneous presence of strong electron correlation leading to local singlet formation, and impurity-metal hybridization, which requires considering the thermodynamic limit. In addition, the Kondo energy scale is very small (typically 1 to 500 K). In low-energy models, one chooses a (small) set of impurity correlated orbitals (e.g., a few $d$ or $f$ orbitals), whose interactions are formally described by a downfolded effective Hamiltonian. Uncertainties, which are large on the Kondo scale, then arise from the choice and construction of these orbitals [Karp2021a], the derivation of the effective interactions (and approximations to treat their complicated frequency dependence [Aryasetiawan2004b]), and the use of double-counting corrections to remove density functional theory (DFT) contributions [Wang2012, Muechler2022b], in lieu of a fully many-body treatment of the local interactions. As a result, although much qualitative progress has been made in describing Kondo physics in different realistic settings [Jacob2009, Surer2012a, Gardonio2013, Dang2016a, Valli2020], the quantitative simulation of Kondo trends across different magnetic impurities, geometries, or even different calculations, without reference to experimental data, is challenging [Surer2012a, Dang2016a, Gardonio2013].
We will instead attempt to solve the ab initio many-electron Schrödinger equation for the Kondo problem without first deriving a low-energy model. This offers the advantage that it is often easier to quantify (and thus converge) errors in the approximate solution of a problem, than the error associated with deriving a model. We build on our work on full cell embedding in the context of dynamical mean-field theories [Zhu2020, Cui2020, Zhu2021c, Cui2022, Cui23-cuprate-doping] to construct a systematically improvable representation of the impurity in its metallic host. Solving the many-body problem in this representation, as we increase the number of orbitals in the parent basis, we eventually obtain the exact non-relativistic, zero-temperature, description of the pure electronic problem. In small molecules [larsson2022chromium], as well as in simple materials (such as organic crystals [yang2014ab]), it has been established that for the quantities of interest, related strategies can reach an accuracy rivalling, or even exceeding, that of experimental measurements. We thus ask, how far can we go with a similar philosophy for a realistic correlated quantum materials problem?
[…]
Results
Numerical strategy.
Our basic plan is to describe the magnetic impurity atom with as complete a basis of electronic orbitals as practical in an impurity-bath embedding setup (Fig. 1a). As the impurity orbital space is increased, it formally converges to the Hilbert space of the material and thus to an exact electronic description (i.e., phonon effects are excluded); for any finite impurity space, the bath approximates the neglected degrees of freedom in the material. While achieving convergence for bulk properties would require including orbitals that span the electronic space of all atoms of the material, here we are focused on observables on the impurity, where convergence of an impurity-centered basis expansion is more rapid. Converging the representation in this way means that we do not need to first downfold the Hamiltonian to derive a model.
More specifically, we study the series of seven $3d$ transition metal impurities (Ti, V, Cr, Mn, Fe, Co, Ni) in bulk Cu, with associated Kondo temperatures that span three orders of magnitude. We started from a Gaussian atomic orbital representation of all atoms: a split-valence double-$\zeta$ polarization basis (def2-SVP) [Weigend2005] for the impurity atoms and a slightly smaller (def2-SV(P)) basis for Cu (which omits the $4f$ basis function). The def2-SVP basis for the impurity atoms contains the $1s2s2p3s3p3d4s4p4d4f5s$ shells (denoted a $5s3p2d1f$ basis, from the shell-count) and thus goes significantly beyond the $3d$ shell considered in a model Hamiltonian treatment. To test the convergence of the results, for a subset of the calculations we also used a larger cc-pVTZ basis [Balabanov2005a] on the impurity atom, corresponding to a $7s6p4d2f1g$ basis. This further improves the electronic treatment of the impurity. We will refer to these as the parent bases and below we demonstrate convergence towards the exact solutions within the parent bases. We estimate the remaining error of the parent basis by the difference between the def2-SVP and cc-pVTZ results.
Starting from DFT-optimized XCu63 structures (X = impurity), we performed periodic DFT calculations in the Gaussian atomic orbital bases with the PBE functional [Perdew96PBE, Sun2020b]. The Gaussian basis functions were then transformed into an orthogonal basis of intrinsic atomic orbitals plus projected atomic orbitals (IAO+PAO) [Knizia2013d]. The impurity IAOs and PAOs are visualized in Fig. 1a. The higher shell orbitals in this picture extend away from the impurity atom onto the neighbours, and because of the close packing of the atoms, start to capture important electronic degrees of freedom of the atoms neighbouring the impurity. For example, the Löwdin population of the spatial extent of the $4s$ orbital of an Fe impurity on the neighbouring Cu atoms is close to $66\%$ in FeCu63. The large basis sets may thus be viewed as forming an “impurity-centered” basis expansion, similar to well-studied local correlation treatments in quantum chemistry.
The impurity-bath embedding Hamiltonian takes the form
$H_{\mathrm{emb}}=\sum_{ij}^{\mathrm{imp}}\tilde{F}_{ij}a_{i}^{\dagger}a_{j}+\frac{1}{2}\sum_{ijkl}^{\mathrm{imp}}(ij|kl)a_{i}^{\dagger}a_{k}^{\dagger}a_{l}a_{j}+\sum_{ip}V_{ip}(a_{i}^{\dagger}c_{p}+c_{p}^{\dagger}a_{i})+\sum_{p}\epsilon_{p}c_{p}^{\dagger}c_{p}$ (1)
where the impurity sum extends over all IAOs and PAOs in the impurity basis, $\{a^{(\dagger)}_{i}\}$ are creation/annihilation operators for impurity states, and $\{c^{(\dagger)}_{p}\}$ are creation/annihilation operators for bath states. The impurity Coulomb interaction matrix $(ij|kl)$ is taken as the bare Coulomb interaction in the IAO+PAO basis, while the impurity one-particle interaction $\tilde{F}_{ij}$ in Eq. 1 is defined as the Hartree-Fock effective Hamiltonian with the local mean-field potential subtracted (this subtraction is diagrammatically exact, so there is no double counting). The impurity-metal hybridization function $\Delta(\omega)$ was obtained at the PBE level. Since we used a Hamiltonian-based impurity solver, we discretized the $3d4s$ valence block of the hybridization function on a non-uniform grid along the real frequency axis [DeVega2015c]
$\Delta_{ij}(\omega)=\sum_{p}\frac{V_{ip}V_{jp}}{\omega-\epsilon_{p}},\leavevmode\nobreak\ ij\in 3d4s$ (2)
where $\{\epsilon_{p}\}$ and $\{V_{ip}\}$ are the bath energies and impurity-bath couplings in Eq. 1. We used 49 bath orbitals to couple to each valence impurity orbital. The total embedding problem thus consisted of 300 electrons in 316 (impurity plus bath) orbitals using the impurity def2-SVP basis, denoted (300e, 316o), and 300 electrons in 353 orbitals using the impurity cc-pVTZ basis, denoted (300e, 353o).
Figure 1: Illustration and validation of computational strategy. (a) All-orbital quantum embedding strategy, where general Coulomb interactions between all valence and high-virtual orbitals of the impurity atom are treated at the quantum many-body level. The percentage of orbital extension onto neighbouring Cu atoms is estimated through the Löwdin population analysis. (b) Experimentally measured Kondo temperature trend ($T_{\mathrm{K}}$ averaged over available experimental values [Daybell1968]) for $3d$ transition metal impurities in bulk Cu, plotted versus the nominal spin ($S$) of isolated impurity atoms. (c) Simulated Kondo resonance and self-energies of the $t_{2g}$ and $e_{g}$ orbitals of the Fe impurity with an active space of (22e, 22o) in the def2-SVP basis. (d) Estimated convergence of the quasiparticle renormalization weights and orbital occupancies of the $3d$ orbitals of Ti, V, and Fe impurities as the DMRG active space increases in the def2-SVP basis. (e) Comparison of quasiparticle renormalization weights and orbital occupancies of $3d$ orbitals of the Fe impurity calculated in the def2-SVP basis versus the cc-pVTZ basis.
To compute and converge the impurity Green’s function, we constructed a series of smaller orbital spaces (active spaces) that are subspaces of the full orbital space of the embedding problem. We then monitored the convergence of the ground-state and dynamical quantities as a function of the active space size. The active spaces were defined to span natural orbitals (eigenvectors of the one-particle density matrix) of an (approximate) ground state of the impurity problem, taking natural orbitals with the largest orbital entropy (a procedure which has been shown to yield highly compact orbital subspaces in model impurity problems [Zgid2012]). We used active spaces of up to (36e, 52o) for the ground-state problem in the impurity def2-SVP basis and (46e, 76o) in the impurity cc-pVTZ basis, and up to (26e, 36o) for the dynamical properties. We computed ground states and associated properties (e.g. density matrices and correlation functions) using ab initio DMRG [Chan2011a, Zhai2021a, Zhai2023a]. For dynamical properties, we used a new zero-temperature ab initio dynamical DMRG (DDMRG) solver to compute the impurity self-energy on the real frequency axis [Ronca2017b, Zhai2023a]. (Benchmarks of the DDMRG solver on the single-impurity Anderson model are shown in Supplementary Note 2).
Convergence towards parent basis and an exact simulation.
We first consider the Fe impurity to illustrate general features of the results, then we discuss numerical convergence. In Fig. 1c we show the orbital-resolved excitation spectra and self-energies of the Fe $3d$ orbitals. As expected, we see sharp Kondo resonance peaks around the Fermi level for both the $t_{2g}$ and $e_{g}$ orbitals, with broad Hubbard shoulder peaks at higher energies. We find that the Fe-$e_{g}$ Kondo peak is narrower and smaller than the Fe-$t_{2g}$ peak from the stronger QP renormalization, which can also be seen from the real part of the real-axis self-energies shown in Fig. 1c. The imaginary part of the real-axis self-energies of both the $t_{2g}$ and $e_{g}$ orbitals approaches zero at the Fermi level (Fig. S5), i.e., $\mathrm{Im}\Sigma(T=0,\omega)\rightarrow 0$ as $\omega\rightarrow 0$. The splitting of the Hubbard peaks is roughly $2.5\sim 3$ eV (in the range of screened interaction for Fe $3d$ orbitals [Nakamura2008a]) and is slightly smaller for $t_{2g}$ orbitals than for $e_{g}$ orbitals.
We then examine the convergence of impurity observables towards the parent basis limit. In Fig. 1d, we show the orbital-resolved QP renormalization weight $Z$ (calculated on real axis as $\Big{[}1-\frac{\partial\Sigma(\omega)}{\partial\omega}\Big{|}_{\omega=0}\Big{]}^{-1}$) and the orbital occupancy $n$ of the $3d$ orbitals, for the Ti, V, and Fe impurities, as a function of active space size with a parent def2-SVP basis. Additional convergence data is in the Supplementary Notes 3 and 4, including the convergence with respect to the DMRG bond dimension (for which the uncertainty is substantially smaller than the uncertainty from the active space for all cases other than Mn). We observe that the QP weight is clearly more challenging to converge. Examining the full series of elements (Supplementary Notes 3 and 4) we conservatively estimate that, with the exception of Mn, $n$ is converged to at least $\sim 0.04$ and $Z$ to at least $\sim 0.05$ of the exact parent basis result. The latter exception arises because it was difficult to fully converge the DMRG calculations in the largest active spaces for Mn.
Beyond these errors, the primary source of deviation from an exact simulation comes from the insufficiency of the parent basis itself. To estimate the error from the parent basis, in Fig. 1e we show the change in $Z$ and $n$ as we significantly increase the parent basis from def2-SVP to cc-pVTZ, for the Fe impurity. We see that both quantities change by less than $\sim 0.03$. Standard quantum chemistry arguments suggest that most quantities converge like the inverse cube of the cardinal number of the basis [helgaker1997basis] (2 in the case of def2-SVP, 3 for cc-pVTZ), which indicates that $Z$ and $n$ are well converged with respect to the representation of the electronic degrees of freedom of the Fe impurity and its close neighbours. Only the contribution from fluctuations involving long-range Coulomb interactions beyond the closest neighbours remains outside of our treatment and this error analysis.
Quantitative Kondo temperature trends and orbital-specific physics.
Having examined the convergence of our numerical results, we now study the trends in Kondo physics across the series of 7 elements. For $3d$ magnetic impurities in bulk Cu, experiments observe an exponential decay in Kondo temperatures moving from nominally low-spin (e.g., Ti) to high-spin (e.g., Mn) impurities [Daybell1968]. This suppression of the Kondo temperature is understood to result from the reduction of the effective exchange coupling $J_{\mathrm{eff}}$, due to Hund’s rule coupling [Georges2013]. Fixing the metallic host, and assuming all other effects are the same for different impurities, this relation can be simplified to
$T_{\mathrm{K}}\propto\exp(-1/J_{\mathrm{eff}})\propto\exp(-S),$ (3)
where $S$ is the spin of the magnetic atom. The estimated experimental Kondo temperatures for the $3d$ magnetic impurities, along with the nominal spin $S$ of the atom is shown in Fig. 1b. Note that the experimental estimates of $T_{\mathrm{K}}$ involve significant interpretation (for a full discussion see [Daybell1968]) and thus a range of characteristic temperatures are obtained depending on the type of measurement and method to extract $T_{\mathrm{K}}$. Nonetheless, Eq. 3 clearly misses quantitative differences between certain magnetic impurities. For example, it does not explain why $T_{\mathrm{K}}$ of Cr (3$\sim$10 K) is lower than $T_{\mathrm{K}}$ of Fe (30 K), if we consider both Cr and Fe atoms to have $S=2$.
To obtain a computational perspective using our methodology, in Fig. 2a we first show the density of states of the Cr, Mn, and Fe impurities. The $t_{2g}$ and $e_{g}$ orbitals of all three impurities yield narrow Kondo resonance peaks around the Fermi level. Since the Kondo temperature $T_{\mathrm{K}}$ is proportional to the width of the Kondo resonance peak, Fig. 2a suggests that the Kondo temperatures of the $t_{2g}$ and $e_{g}$ orbitals follow different trends: $T_{\mathrm{K}}(\mathrm{Cr})<T_{\mathrm{K}}(\mathrm{Mn})<T_{\mathrm{K}}(\mathrm{Fe})$ for the $t_{2g}$ orbitals and $T_{\mathrm{K}}(\mathrm{Mn})<T_{\mathrm{K}}(\mathrm{Fe})\approx T_{\mathrm{K}}(\mathrm{Cr})$ for the $e_{g}$ orbitals. This is corroborated by the behaviour of the real-frequency self-energies (Fig. S5), and illustrates that the trend across the elements is intertwined with orbital-specific physics.
Figure 2: Trends in impurity DOS, quasiparticle renormalization, spin moment, and Kondo temperature. All results are in the def2-SVP basis with ground-state active space of (36e, 52o) and Green’s function active space of (22e, 22o). (a) Comparison of DOS of $3d$ orbitals of Cr, Mn, and Fe impurities. (b) Top: Orbital-resolved quasiparticle renormalization weights $Z$. Bottom: Local spin moments of impurity atoms compared to nominal spins of isolated atoms and experimental values. (c) Kondo temperatures from all-orbital DMFT predictions compared to experimental values [Daybell1968]. The experimental trend lines are provided as a visual aid. Symbols for each type of measurements are: $\rho$: resistivity, $\chi$: susceptibility, $C$: specific heat, $S$: thermoelectric power.
For a more quantitative estimate, we extracted Kondo temperatures for the $t_{2g}$ and $e_{g}$ orbitals for all seven $3d$ magnetic impurities using the relation from Hewson’s renormalized perturbation theory of the Anderson impurity model [Hewson1993]
$T_{\mathrm{K}}=-\frac{\pi}{4}Z\cdot\mathrm{Im}\Delta(0).$ (4)
Here $\Delta(0)$ is the hybridization function at the Fermi level and $Z$ is the QP renormalization weight on the real axis (Fig. S4 and Table S4). $\Delta(0)$ only varies slightly across the series (from $-0.59$ eV (Ti) to $-0.43$ eV (Fe) for the $t_{2g}$ orbital and from $-0.64$ eV (Ti) to $-0.38$ eV (Ni) for the $e_{g}$ orbital). On the other hand, we see a dramatic reduction of the QP renormalization weight $Z$ moving from low-spin to high-spin impurities for both the $t_{2g}$ and $e_{g}$ orbitals, as seen in Fig. 2b, and the trend is orbital-dependent. For the $t_{2g}$ orbitals, Cr has the lowest QP weight, at $Z=0.017$, and the trend for QP weight follows: Cr $<$ Mn $<$ Fe $<$ Co $\approx$ V $<$ Ni $<$ Ti. For the $e_{g}$ orbitals, Mn has the strongest QP renormalization with $Z=0.006$ and the trend for QP weight follows: Mn $<$ Fe $<$ Cr $<$ Co $<$ V $<$ Ni $<$ Ti.
From Eq. 4, we obtain $T_{\mathrm{K}}(t_{2g})=73\leavevmode\nobreak\ \mathrm{K}$ and $T_{\mathrm{K}}(e_{g})=167\leavevmode\nobreak\ \mathrm{K}$ for Cr, and $T_{\mathrm{K}}(t_{2g})=565\leavevmode\nobreak\ \mathrm{K}$ and $T_{\mathrm{K}}(e_{g})=101\leavevmode\nobreak\ \mathrm{K}$ for Fe. The lower of the two Kondo temperatures is the relevant one for the measurements (see e.g., Ref. [daybell1967observation] for resistivity), thus we obtain characteristic Kondo temperatures of Cr (73 K) and Fe (101 K). This indeed reproduces the subtle experimental trend, which we see is driven by electron correlation in different orbitals in Cr ($t_{2g}$) and Fe ($e_{g}$) respectively.
More insights can be obtained from the computed local ground-state properties of the impurity, including the $t_{2g}$ and $e_{g}$ orbital occupancies across the impurity elements (Table S5) and the effective spin (extracted from $\langle S^{2}\rangle$) in Fig. 2b. Charge transfer takes place to the impurity atom in all systems, giving negatively charged $3d$ shells compared to the isolated impurities. For instance, we find Fe in bulk Cu to be on average in a $d^{7}$ occupancy ($n_{\mathrm{Fe}}(3d)=6.92$), in excellent agreement with recent experimental estimates [Joly2017]. The strong QP renormalization in Fig. 2b tracks the partially-filled nature of the impurity $3d$ orbitals. For example, we obtain $n(t_{2g})=0.99$ in Cr and $n(e_{g})=1.03$ in Mn (i.e., very close to half filling), which correlates well with their very small QP weights.
As a result of the charge transfer, the local spin moments of the magnetic impurities differ from their isolated-atom values (Fig. 2b). For Mn, Fe, Co, and Ni, we predict $S=1.99,1.42,0.95,0.48$, about 0.5 lower than the isolated-atom spins. These $S$ values agree well with the available experimental data for Mn and Fe [Daybell1968, Haen1976, Joly2017]. Cr has the highest spin moment ($S=2.23$) among all $3d$ transition metal impurities, which deviates from the common understanding of this Kondo series [Nevidomskyy2009], but is also supported by experimental measurements that suggest Cr is a high-spin impurity ($S=5/2$[Haen1976, Sacramento1990]. From Eq. 3, the large moment of Cr in the Cu host suggests that the observed suppression of $T_{\mathrm{K}}$ relative to that of Fe arises from the same physical principles. However, it is also clear that the trends in $T_{\mathrm{K}}$ are not solely captured by the effective $S$ (see Supplementary Note 8 for orbital-resolved spin correlations).
Our predicted Kondo temperatures for all the studied elements are shown in Fig. 2c. We capture the exponential decay in the Kondo temperature moving from low-spin to high-spin impurities, in good agreement with the experimental estimates. The Kondo energy scales of Mn, Fe, Co, and Ni are found to be $e_{g}$-driven, while they are $t_{2g}$-driven for Cr, V, and Ti. With the exception of Cr and Mn, our predicted $T_{\mathrm{K}}$ is within a factor of 2$\sim$3 of at least one of the experimental estimates of $T_{\mathrm{K}}$. As our calculations do not include long-range Coulomb fluctuations (beyond those captured by the basis extending to neighbouring atoms), this suggests that they do not contribute significantly at this level of accuracy. On the other hand, Mn, which has the lowest $T_{\mathrm{K}}$, is a system where we could not achieve numerical convergence of our results to the parent basis. This is the likely origin of the larger quantitative discrepancy.
Figure 3: Comparison between model and ab initio all-orbital simulations. (a) Orbital-resolved quasiparticle renormalization weights. (b) Predictions of Kondo temperatures against experimental values.
Comparison to low-energy models.
Finally, to place the fidelity of our results in perspective, we check to see if the above predictions are easily captured within a standard application of the downfolded model approach. For this, we employ a five-orbital Anderson impurity model using the Kanamori Hamiltonian [Kanamori1963a, Georges2013] with screened Coulomb interaction and exchange parameters $U$ and $J$ taken from constrained random phase approximation (cRPA) calculations [Sasioglu2011]. (Note these parameters are also very close to previous parameters used in other model Kondo studies, e.g., for Co in Cu [Surer2012a, Valli2020]). We employed the same valence hybridization, bath discretization, and active-space DMRG solvers described above, and further details are given in Supplementary Note 9. Fig. 3 shows the estimated Kondo temperatures and quasiparticle renormalization obtained from this model. We find that the quasiparticle renormalization weights of most magnetic impurities are significantly overestimated compared to our all-orbital simulations. As a result, the predicted $T_{\mathrm{K}}$’s do not even show a clear exponential trend, with a large overestimation of $T_{\mathrm{K}}$. In addition, the relative $T_{\mathrm{K}}$ of Cr and Fe, one of the more interesting and subtle trends discussed above, is not reproduced in the model calculation. The quantitative improvement in the Kondo temperatures moving from the model calculations can be quite large: as much as an order of magnitude, or even more.
Discussion.
We have demonstrated the predictive power of ab initio quantum many-body simulations in the prototypical Kondo physics materials problem, namely, that of $3d$ transition metal impurities in bulk copper. By converging the material description and many-body treatment, we could reach an accuracy for the Kondo temperature that captures the subtle experimental trends across the $3d$ transition metal series.
In contrast to previous approaches, we achieved this accuracy by performing the quantum many-body simulations of the bare electronic problem with all orbitals, rather than within a downfolded model. This allows for a straightforward control of many aspects of converging our results towards an exact description of the phenomena. Within the zero-temperature electronic picture of our work, the physics of very long-range Coulomb fluctuations remain as an uncontrolled uncertainty, although these effects appear to be small on the scale of our results. However, our computational framework does not lie in opposition to model approaches. In particular, now that we can establish faithful physics within our framework, we can seek to extract traditional low-energy models from it. Moving forwards, the techniques used here to simulate Kondo physics are broadly applicable to more complicated correlated electron materials. The accuracy achieved here suggests that with continuing advances, we may move beyond qualitative descriptions of correlated electron phenomena, towards the quantitative simulation of observables directly measured by experiments.
Data availability
Data used in this work are available in the main text and supplementary materials.
Code availability
The fcDMFT code is available at https://github.com/ZhuGroup-Yale/fcdmft. The libDMET code is available at
https://github.com/gkclab/libdmet_preview. The block2 code is available at https://github.com/block-hczhai/block2-preview. PySCF is available at https://github.com/pyscf/pyscf.
Acknowledgements
In the initial phase of this project performed at Caltech, T.Z. was supported by the Air Force Office of Scientific Research through the MURI program, Grant No. FA9550-18-1-0095. The development of integral infrastructure by Z.-H.C. was supported by the US Department of Energy, Office of Science, Basic Energy Sciences, through award DE-SC0018140. L.P.’s continuation of the work by T.Z. was supported by the Center for Molecular Magnetic Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Grant No. DE-SC0019330. T.Z. also acknowledges support from the Air Force Office of Scientific Research under award number FA9550-24-1-0096 for the work conducted at Yale. H.Z. (DMRG solver) was supported by the US National Science Foundation, through Grant No. CHE-2102505. Calculations were conducted in the Resnick High Performance Computing Center, supported by the Resnick Sustainability Institute at Caltech, as well as in the Yale Center for Research Computing.
Author Contributions
T.Z. and G.K.-L.C. designed the initial project and wrote the manuscript. T.Z., L.P., and H.Z. developed all-orbital DMFT method and code. T.Z., L.P., and Z.-H.C. performed ab initio calculations of Kondo systems. T.Z. carried out the model Hamiltonian calculations. H.Z. developed the code of DMRG impurity solver. G.K.-L.C. supervised the project. All authors contribute to the discussion of the results as well as the writing and editing of the manuscript.
Competing interests
G.K.-L.C. is a part owner of QSimulate.
Additional information
Methods, Figures S1-S6, Tables S1-S8.
[…]
Supplementary Materials for:
Towards an exact electronic quantum many-body treatment of Kondo correlation in magnetic impurities
Supplementary Note 1 Computational methods
We generated initial XCu63 (X = Ti, V, Cr, Mn, Fe, Co, Ni) structures by replacing one Cu atom in a $4\times 4\times 4$ supercell of bulk Cu with an impurity atom. We carried out DFT geometry relaxations for all XCu63 structures with the PBE functional and projector-augmented-wave (PAW) basis using the Vienna Ab initio Simulation Package (VASP) [Kresse1996, Kresse1999]. The calculations were performed with a plane wave cutoff of 400 eV and a $\Gamma$-centered $3\times 3\times 3$ $\mathbf{k}$-mesh. The forces on each atom were converged to less than 0.01 eV/Å. With DFT-optimized structures, we performed single-point DFT calculations with the PBE functional in a periodic Gaussian basis set using the PySCF quantum chemistry software package [Sun2020b]. We employed a split-valence double-$\zeta$ polarization basis (def2-SVP) for all impurity atoms and a split-valence double-$\zeta$ basis (def2-SV(P)) for Cu [Weigend2005]. The correlation-consistent polarized triple-$\zeta$ basis (cc-pVTZ) was also employed for the Fe impurity to test the parent basis convergence [Balabanov2005a]. The libDMET code [Cui2020, Cui2022] was used to transform mean-field Fock matrix, density matrix, and electron repulsion integrals to the intrinsic atomic orbital plus projected atomic orbital (IAO+PAO) basis [Knizia2013d]. $3d4s$ atomic orbitals of all metal atoms were used as the predefined valence (minimal) orbitals in the IAO+PAO construction.
We incorporate the full impurity atom into the embedding problem. In the IAO+PAO basis, the $3d4s4p4d4f5s$ (def2-SVP) and $3d4s4p4d4f5s5p5d5f5g6s6p6d7s7p$ (cc-pVTZ) orbitals of the impurity atoms were treated in the many-body impurity solvers, while the $1s2s2p3s3p$ orbitals were frozen at the mean-field level and left out of the embedding problem. The fcDMFT code [Zhu2019, Zhu2020, Zhu2021c] was used to perform the all-orbital DMFT calculations. We employed the bare Coulomb interaction $(ij|kl)$ within the impurity as the two-particle interaction matrix in the embedding Hamiltonian in Eq. 1. The one-particle impurity interaction matrix in Eq. 1 is defined as
$\tilde{F}_{ij}=F^{\mathrm{imp}}_{ij}-\sum_{kl\in\mathrm{imp}}\gamma^{\mathrm{imp}}_{kl}[(ij|lk)-\frac{1}{2}(ik|lj)],$ (S1)
where $F^{\mathrm{imp}}$ is the impurity Fock matrix computed by Hartree-Fock using the PBE density, and $\gamma^{\mathrm{imp}}$ is the impurity block of PBE density matrix. This definition ensures that there is no double counting in the impurity Hamiltonian.
The hybridization function for each impurity atom was obtained at the PBE level using a $4\times 4\times 4$ ${\mathbf{k}}$-point sampling of the XCu63 supercell:
$\Delta(\omega+i\delta)=\omega+i\delta-F_{\mathrm{imp}}-G_{\mathrm{imp}}^{-1}(\omega+i\delta),$ (S2)
where $\delta$ is the broadening factor taken to be $\delta=0.01$ Ha. Since Hamiltonian-based impurity solvers (that require a bath) were employed, we discretized the $3d4s$ block of the hybridization function on a non-uniform grid along the real frequency axis using a pole-merging direct discretization method [DeVega2015c] to obtain bath energies $\{\epsilon_{p}\}$ and impurity-bath couplings $\{V_{ip}\}$ in Eq. 1, and 49 bath orbitals were coupled to each of the $3d4s$ valence orbitals. Among these 49 bath orbitals, 5 orbitals were placed within a $\mu\pm 0.027$ eV energy window, 18 orbitals were placed within the $[\mu\pm 0.027,\mu\pm 0.6]$ eV energy window, and the remaining 26 orbitals were placed at a higher energy window of $[\mu\pm 0.6,\mu\pm 8.0]$ eV. In total, the embedding problem contains 22 (impurity) + 294 (bath) = 316 orbitals in the def2-SVP impurity basis calculations. The same bath discretization procedure was applied to the cc-pVTZ impurity basis calculation of the Fe impurity, resulting in an embedding problem of 59 (impurity) + 294 (bath) = 353 orbitals.
To solve the embedding Hamiltonian, we first performed a Hartree-Fock calculation with the chemical potential fixed at the supercell DFT value, to define the number of electrons in the embedding problem. We note that the value of the chemical potential is often chosen to tune the orbital fillings, but we found that the current strategy gave excellent agreement with experimental occupancies of Mn and Fe. Following the HF solution, we carried out a configuration interaction with singles and doubles (CISD) calculation and computed natural orbitals by diagonalizing the CISD one-particle density matrix. A (36e, 52o) natural-orbital active space was then derived in the def2-SVP impurity basis calculations ((36e, 46o) for Ni, (46e, 76o) in the cc-pVTZ impurity basis calculation of Fe), where all kept natural orbitals have eigenvalues $n_{i}$ that satisfy $\mathrm{min}(n_{i},2-n_{i})>10^{-7}$. An ab initio quantum chemistry DMRG calculation was then conducted on this active space to obtain ground-state properties including one-particle and two-particle density matrices and spin-spin correlation functions. The DMRG calculation was done with a bond dimension of $M=3500$ ($M=4000$ for Mn and Fe) using the block2 code [Zhai2021a, Zhai2023a], where the discarded weight was below $2\times 10^{-5}$ in all ground-state DMRG calculations. We further derived a series of smaller natural-orbital active spaces from the DMRG density matrix (see Supplementary Note 3 for details). The dynamical DMRG (DDMRG) [Ronca2017b] calculation was carried out for the smaller active spaces with a bond dimension up to $M=1500$ along the real axis at zero temperature. Larger bond dimensions of up to $M=4000$ were employed for strongly correlated sites to ensure that the discarded weight in the DDMRG calculations was below 0.02. To accommodate the non-uniform bath discretization, we used a broadening factor of $\eta=0.02$ eV within the $\mu\pm 0.25$ eV energy window, $\eta=0.05$ eV within the $[\mu\pm 0.25,\mu\pm 0.5]$ eV energy window, $\eta=0.2$ eV within the $[\mu\pm 0.5,\mu\pm 2.0]$ eV energy window, and $\eta=0.6$ eV within the $[\mu\pm 2.0,\mu\pm 5.0]$ eV energy window in the DDMRG calculations.
From the DDMRG calculations, we extracted the self-energy of the active space $\Sigma_{\mathrm{DMRG,act}}(\omega)$ as
$\Sigma_{\mathrm{DMRG,act}}(\omega)=G_{\mathrm{DFT,act}}^{-1}(\omega)-G_{\mathrm{DMRG,act}}^{-1}(\omega),$ (S3)
where $G_{\mathrm{DFT,act}}$ was calculated from the effective DFT Hamiltonian rotated to the active space. The active-space self-energy was then rotated back to the full embedding space
$\Sigma_{\mathrm{DMRG,emb}}(\omega)=C^{\mathrm{emb,act}}\Sigma_{\mathrm{DMRG,act}}(\omega)(C^{\mathrm{emb,act}})^{\dagger},$ (S4)
where $C^{\mathrm{emb,act}}=C^{\mathrm{CISD}}C^{\mathrm{DMRG}}$. Here, $C^{\mathrm{CISD}}$ represents the rotation matrix from the full embedding space to the natural-orbital active space derived from CISD calculations, while $C^{\mathrm{DMRG}}$ represents the rotation matrix from the CISD active space to the natural-orbital active space derived from ground-state DMRG calculations. Finally, the local Green’s function of the impurity was calculated from the Dyson’s equation
$G_{\mathrm{loc}}(\omega)=[G_{\mathrm{DFT,emb}}^{-1}(\omega)-\Sigma_{\mathrm{DMRG,emb}}(\omega)]^{-1}.$ (S5)
Supplementary Note 2 Benchmark of DDMRG solver on Anderson impurity model
We benchmark the accuracy of the active-space DMRG and dynamical DMRG (DDMRG) solvers on a single-impurity Anderson model (SIAM), the fundamental model of Kondo physics, where high-accuracy numerical results are available (e.g., from NRG [Bulla1998] or DMRG [Nishimoto2004a]). The Hamiltonian of the SIAM is
$H=\sum_{\sigma}\epsilon_{f}f^{\dagger}_{\sigma}f_{\sigma}+Uf^{\dagger}_{\uparrow}f_{\uparrow}f^{\dagger}_{\downarrow}f_{\downarrow}+\sum_{k\sigma}\epsilon_{k}c^{\dagger}_{k\sigma}c_{k\sigma}+\sum_{k\sigma}V_{k}(f^{\dagger}_{\sigma}c_{k\sigma}+c^{\dagger}_{k\sigma}f_{\sigma}),$ (S6)
where $f^{(\dagger)}_{\sigma}$ are creation/annihilation operators for impurity states with spin $\sigma$ and energy $\epsilon_{f}$, $c^{(\dagger)}_{k\sigma}$ are creation/annihilation operators for band states with spin $\sigma$ and energy $\epsilon_{k}$, $U$ is the impurity on-site Coulomb interaction, and $V_{k}$ are the $k$-dependent coupling between impurity and band states. We followed Ref. [Bulla1998] and employed the flat-band hybridization function:
$\mathrm{Im}\Delta(\omega+i0^{+})=-0.015D,\leavevmode\nobreak\ |\omega|<D$ (S7)
where $2D$ is the conduction electron bandwidth. For convenience, we set $D=1$. The flat-band hybridization on a logarithm grid was then discretized along the real axis to obtain $\{\epsilon_{k},V_{k}\}$:
$-\mathrm{Im}\Delta(\omega+i0^{+})=\sum_{k}\frac{V_{k}^{2}}{\omega-\epsilon_{k}}.$ (S8)
We coupled 39 bath orbitals to the single impurity orbital. Furthermore, only the half-filling case was considered, which means $\epsilon_{f}=-\frac{1}{2}U$. We tested our impurity solver on three interaction strengths: $U=0.1,0.2,0.3$.
We first solved the SIAM Hamiltonian using the Hartree-Fock approximation, followed by a configuration interaction with single and double excitations (CISD) calculation. We then derived a natural-orbital-based active space by diagonalizing the CISD one-particle density matrix. We denote the $n$-electron $n$-orbital active space as ($n$e, $n$o). We then solved the active space using ground-state DMRG with bond dimension $M=1200$ and then dynamic DMRG with bond dimension $M=800$ [Zhai2021a, Zhai2023a]. To understand the accuracy of this active-space DMRG solver, we also solved the full 40-orbital SIAM problem using DMRG/DDMRG with the same bond dimensions.
Figure S1: Impurity density of states and self-energy in the single-impurity Anderson model calculated by full DMRG and active-space DMRG. $U=0.2$. (a) (4e, 4o) active space. (b) (8e, 8o) active space. (c) (12e, 12o) active space.
We first present the impurity density of states (DOS) and self-energy at $U=0.2$ in Fig. S1. It shows a sharp Kondo resonance peak and two broad Hubbard bands in the full DMRG calculation, which agrees quantitatively with previous NRG results [Bulla1998]. Compared to the full (40e, 40o) DMRG results (Fig. S1a), the (4e, 4o) active-space result shows multiple spurious oscillations in the DOS and self-energy. However, we point out that even with a very small (4e, 4o) active space, the impurity solver correctly predicts the shape of the Kondo resonance around the Fermi level, which is also confirmed in the self-energy comparison around $\omega=0$. We find that the accuracy improves rapidly when the active space is increased from (4e, 4o) to (8e, 8o) and (12e, 12o). At the (12e, 12o) level, the DOS and self-energy from active-space DMRG are almost indistinguishable from the full DMRG results.
Figure S2: Quasiparticle renormalization weight $Z$ for increasing sizes of active space in $U=0.1,0.2,0.3$ SIAM models. The full (40e, 40o) DMRG calculated $Z$ are shown as dashed lines.
In Fig. S2, we further compare the quasiparticle renormalization factors $Z$ defined as
$Z=\Big{[}1-\frac{\partial\Sigma(\omega)}{\partial\omega}\Big{|}_{\omega=0}\Big{]}^{-1}.$ (S9)
We find that at moderate interaction strength ($U=0.1$), the active-space DMRG is already accurate at the (12e, 12o) level, predicting $Z=0.22$ vs. the full DMRG result $Z=0.20$. At stronger interaction strength ($U=0.2$), the (12e, 12o) active-space DMRG predicts $Z=0.035$, an overestimation compared to the full DMRG result $Z=0.018$. In summary, for stronger interactions ($U=0.2$ and $U=0.3$), a larger active space than (12e, 12o) is needed to achieve a completely quantitative description of the quasiparticle renormalization. Nevertheless, at all active space sizes, the active-space DMRG solver correctly predicts the exponential decay in $Z$ as $U$ increases from 0.1 to 0.3. We point out that it is difficult to directly infer the accuracy of the active-space DMRG in the ab initio Kondo simulations from the SIAM calculations alone because of the existence of multiple $3d$ orbitals and the inclusion of all other impurity valence and virtual orbitals into the embedding problem. However, the convergence of the quantities in the SIAM calculation with the different components of the calculation procedure (which are the same in the SIAM and ab initio calculations) supports the convergence checks that we discuss in the main text.
Supplementary Note 3 Convergence of impurity observables towards parent basis limit
Table S1: Occupancies ($n$) and quasiparticle renormalization weights ($Z$) of impurity $t_{2g}$ and $e_{g}$ orbitals obtained from different DMRG/DDMRG active-space simulations. All results are in the def2-SVP basis unless specified.
Impurity Active space $n(t_{2g})$ $n(e_{g})$ $Z(t_{2g})$ $Z(e_{g})$
Ti (10e, 10o) 0.59 0.26 0.64 0.59
(18e, 18o) 0.60 0.30 0.59 0.62
(22e, 22o) 0.60 0.32 0.58 0.58
(30e, 29o) 0.61 0.30 0.60 0.59
(36e, 52o) 0.62 0.32
V (10e, 10o) 0.99 0.19 0.27 0.45
(18e, 18o) 0.93 0.37 0.20 0.47
(22e, 22o) 0.93 0.39 0.19 0.37
(28e, 30o) 0.93 0.38 0.19 0.37
(36e, 52o) 0.95 0.41
Cr (10e, 10o) 0.91 0.82 0.12 0.12
(14e, 16o) 1.10 0.61 0.092 0.16
(22e, 22o) 0.97 0.94 0.017 0.038
(28e, 28o) 0.99 0.93 0.023 0.047
(36e, 52o) 0.99 0.96
Mn (10e, 10o) 1.06 1.07 0.045 0.011
(16e, 16o) 1.16 1.00 0.068 0.018
(22e, 22o) 1.17 1.03 0.071 0.0058
(36e, 52o) 1.22 1.03
Fe (10e, 10o) 1.46 1.14 0.27 0.12
(def2-SVP) (16e, 16o) 1.59 1.03 0.23 0.021
(22e, 22o) 1.58 1.06 0.15 0.028
(28e, 28o) 1.59 1.04 0.15 0.026
(36e, 52o) 1.58 1.08
Co (10e, 10o) 1.85 1.16 0.29 0.11
(16e, 16o) 1.90 1.09 0.37 0.060
(22e, 27o) 1.80 1.26 0.19 0.11
(26e, 36o) 1.79 1.28 0.20 0.11
(36e, 52o) 1.77 1.32
Ni (10e, 10o) 1.77 1.77 0.50 0.47
(16e, 16o) 1.74 1.85 0.48 0.47
(22e, 22o) 1.80 1.80 0.42 0.40
(28e, 28o) 1.78 1.81 0.40 0.38
(36e, 46o) 1.80 1.82
Fe (26e, 29o) 1.54 1.08 0.19 0.039
(cc-pVTZ) (46e, 76o) 1.56 1.08
Figure S3: Convergence of the quasiparticle renormalization weights (left) and orbital occupancies (right) of the $3d$ orbitals of all impurities as the DMRG active space increases.
In this section, we present detailed benchmarks of the convergence of local impurity observables in the active-space DMRG calculations towards the parent basis limit. By diagonalizing the one-particle density matrix obtained from ground-state DMRG calculations within the (36e, 52o) active space ((36e, 46o) for Ni) in the def2-SVP basis, we further derived a series of smaller active spaces for performing the more time-consuming dynamical DMRG (DDMRG) calculations. The tested active spaces are listed in Table S1, where the corresponding orbital occupancies $n$ and quasiparticle (QP) renormalization weights $Z$ of impurity $3d$ orbitals are shown in Table S1 and Fig. S3. We find that, with the exception of Mn (where we could not converge the DDMRG calculation with respect to bond dimension for an active space larger than (22e, 22o)), $Z$ and $n$ are converged with small errors for all impurities at an active-space size of (22e, 22o) ((22e, 27o) for Co).
Supplementary Note 4 Many-body solver error estimates
We next present an estimate of the errors from two of the approximations in our calculations: active space size and DMRG bond dimension, which are the main many-body solver errors with respect to the exact parent-basis limit (in this case, def2-SVP). First, we analyze the errors in the predicted impurity observables due to the use of active spaces. As shown in Table S2, for orbital occupancies, we estimate the remaining errors relative to the parent-basis limit by comparing the largest and second largest ground-state (GS) active-space calculations (assuming (36e, 52o) is a sufficiently large active space). For example, for Fe, we obtain $\Delta n(t_{2g})=0.010$ and $\Delta n(e_{g})=0.045$ by comparing the 52o (largest GS) and 28o (second largest GS) results. Using the maximum $\Delta n$ across all impurities (except Mn), we estimate that the active space associated orbital occupancy error to be at most 0.045. For estimating the QP weight errors, we adopt a two-step procedure: (1) Compare the largest and second largest Green’s function (GF) active-space calculations. For Fe, we obtain $\Delta Z(t_{2g})=0.009$ and $\Delta n(e_{g})=0.002$ by taking the difference between the 28o (largest GF) and 22o (second largest GF) results. (2) Estimate the QP weight errors between the largest and second largest GS active spaces by assuming a linear relationship between $\Delta n$ and $\Delta Z$:
[…]
Following this procedure, we estimate the $Z$ errors for Fe to be $\Delta Z(t_{2g})=0.008$ and $\Delta n(e_{g})=0.005$. Furthermore, using the maximum $\Delta Z$ across all impurities (except Mn), we estimate the active space associated QP weight error to be at most 0.048.
Table S2: Active-space error analysis of orbital occupancies ($n$) and quasiparticle renormalization weights ($Z$) of impurity $3d$ impurity orbitals. The “est.” values are estimated according to Eq. S10.
Impurity Comparison btw. active spaces $\Delta n(t_{2g})$ $\Delta n(e_{g})$ $\Delta Z(t_{2g})$ $\Delta Z(e_{g})$
Ti 22o vs. 29o 0.011 0.020 0.015 0.008
29o vs. 52o 0.013 0.015 0.018 (est.) 0.006 (est.)
V 22o vs. 30o 0.002 0.011 0.007 0.001
30o vs. 52o 0.014 0.022 0.048 (est.) 0.001 (est.)
Cr 22o vs. 28o 0.015 0.018 0.005 0.010
28o vs. 52o 0.001 0.030 0.000 (est.) 0.016 (est.)
Fe 22o vs. 28o 0.011 0.019 0.009 0.002
28o vs. 52o 0.010 0.045 0.008 (est.) 0.005 (est.)
Co 22o vs. 36o 0.009 0.022 0.006 0.005
36o vs. 52o 0.013 0.038 0.010 (est.) 0.008 (est.)
Ni 22o vs. 28o 0.011 0.013 0.020 0.024
28o vs. 46o 0.015 0.007 0.029 (est.) 0.013 (est.)
Table S3: Error analysis of quasiparticle renormalization weights of impurity $3d$ orbitals against DDMRG maximum bond dimensions and discarded weights. The “extrap.” values are obtained through two-point linear extrapolation against discarded weights.
Impurity Max. bond dimension Max. discarded weight $Z(t_{2g})$ $Z(e_{g})$
V 2500 $1.52\times 10^{-2}$ 0.1893 0.3742
(28e, 30o) 4000 $9.41\times 10^{-3}$ 0.1867 0.3736
extrap. 0 0.1825 0.3726
||||$Z$($M=4000$)--$Z$(extrap.)|||| $4.2\times 10^{-3}$ $1.0\times 10^{-3}$
Cr 2500 $2.24\times 10^{-2}$ 0.0202 0.0473
(28e, 28o) 3500 $1.47\times 10^{-2}$ 0.0226 0.0474
extrap. 0 0.0273 0.0476
||||$Z$($M=3500$)--$Z$(extrap.)|||| $4.7\times 10^{-3}$ $1.9\times 10^{-4}$
Fe 2500 $5.28\times 10^{-3}$ 0.1538 0.0253
(28e, 28o) 3500 $2.97\times 10^{-3}$ 0.1538 0.0255
extrap. 0 0.1539 0.0258
||||$Z$($M=3500$)--$Z$(extrap.)|||| $7.3\times 10^{-5}$ $2.9\times 10^{-4}$
Co 1500 $3.46\times 10^{-2}$ 0.2011 0.1088
(26e, 36o) 2500 $1.29\times 10^{-3}$ 0.2003 0.1071
extrap. 0 0.1999 0.1062
||||$Z$($M=2500$)--$Z$(extrap.)|||| $4.5\times 10^{-4}$ $9.9\times 10^{-4}$
We then analyze the numerical errors due to finite bond dimension in DMRG calculations. We find that the DMRG predicted orbital occupancies are very well converged with respect to the bond dimension ($M$) in all impurities. For example, for Cr, the orbital occupancy differences between $M=3000$ and $M=4000$ calculations are negligibly small: $\Delta n(t_{2g})=2\times 10^{-4}$ and $\Delta n(e_{g})=1\times 10^{-4}$.
We then focus on the DMRG bond dimension associated QP weight errors. In Table S3, we present DDMRG $Z$ values and discarded weights at various bond dimensions for V, Cr, Fe, and Co, in their largest GF active-space calculations. We also performed a two-point linear extrapolation against the discarded weights to estimate the $Z$ values at infinite DDMRG bond dimension. By comparing the largest $M$ results against the extrapolated values, we find that the largest bond dimension associated $Z$ error is around $5\times 10^{-3}$ (Cr), an order of magnitude smaller than the largest active space associated $Z$ error.
In summary, combining the maximum errors from the active space and bond dimension error analysis, we conservatively estimate that our predicted $n$ values are converged to at least $\sim$ 0.04 and $Z$ values are converged to at least $\sim$ 0.05 compared to the exact parent basis result.
Supplementary Note 5 DFT hybridization functions
We present DFT-calculated real-axis hybridization functions of magnetic impurity atoms in bulk Cu in Fig. S4. It is observed that the $t_{2g}$ hybridization has a greater magnitude than the $e_{g}$ hybridization in all impurities, especially in the range of [3, 6] eV. Meanwhile, we find that the magnitudes of both $t_{2g}$ and $e_{g}$ hybridization functions become smaller from Ti to Ni.
Figure S4: Real-axis hybridization function of $3d$ orbitals of magnetic impurities in bulk Cu calculated by DFT with the PBE functional.
Figure S5: Real-axis self-energies of Cr, Mn, and Fe impurities from all-orbital DMFT. (a) Real part of self-energies of $t_{2g}$ orbitals. (b) Real part of self-energies of $e_{g}$ orbitals. (c) Imaginary part of self-energies of $t_{2g}$ orbitals. (d) Imaginary part of self-energies of $e_{g}$ orbitals.
Supplementary Note 6 Self-energies of magnetic impurity orbitals
We show real-axis self-energies of Cr, Mn, and Fe impurities in bulk Cu calculated by all-orbital DMFT in Fig. S5. When $\omega$ approaches the Fermi level, the imaginary part of the self-energies of the $3d$ orbitals approaches zero, while the real part of the self-energies of the $3d$ orbitals changes linearly with respect to the frequency, which agrees with the expected Fermi liquid behavior.
Supplementary Note 7 Summary of ground-state and spectral properties
The ground-state and spectral properties of magnetic impurities calculated by all-orbital DMFT are summarized in Table S4 and Table S5. In Table S5, we include the natural orbital occupancies of $t_{2g}$ and $e_{g}$ symmetry derived from the (36e, 52o) active-space DMRG density matrix. We choose to show the most fractional occupancies with $n_{\mathrm{nat}}>1$. We observe that the natural occupancies in Table S5 correlate well with the quasiparticle renormalization $Z$ in Table S4, i.e., more fractionally occupied orbitals are found to possess stronger quasiparticle renormalization.
Table S4: Hybridization function values at the Fermi level $\Delta(0)$, quasiparticle renormalization weights $Z$, and Kondo temperatures $T_{\mathrm{K}}$ of magnetic impurities in bulk Cu computed by all-orbital DMFT. All results are in the def2-SVP basis with Green’s function active space of (22e, 22o) ((22e, 27o) for Co).
Impurity $\Delta(0,t_{2g})$ (eV) $\Delta(0,e_{g})$ (eV) $Z(t_{2g})$ $Z(e_{g})$ $T_{\mathrm{K}}(t_{2g})$ (K) $T_{\mathrm{K}}(e_{g})$ (K)
Ti $-0.59$ $-0.64$ 0.58 0.58 3144 3412
V $-0.52$ $-0.56$ 0.19 0.37 924 1895
Cr $-0.47$ $-0.48$ 0.017 0.038 73 167
Mn $-0.44$ $-0.43$ 0.071 0.0058 287 23
Fe $-0.43$ $-0.40$ 0.15 0.028 565 101
Co $-0.43$ $-0.38$ 0.19 0.11 761 396
Ni $-0.44$ $-0.38$ 0.42 0.40 1709 1374
Table S5: Orbital occupancies $n$, natural orbital occupancies $n_{\mathrm{nat}}$, and spin moments $S$ of magnetic impurities in bulk Cu computed by all-orbital DMFT. All results are in the def2-SVP basis with ground-state active space of (36e, 52o) ((36e, 46o) for Ni).
Impurity $n(t_{2g})$ $n(e_{g})$ $n_{\mathrm{nat}}(t_{2g})$ $n_{\mathrm{nat}}(e_{g})$ $n(\mathrm{3d})$ $S$
Ti 0.62 0.32 1.84 1.93 2.50 0.93
V 0.95 0.40 1.57 1.86 3.65 1.36
Cr 0.99 0.96 1.24 1.21 4.87 2.23
Mn 1.22 1.03 1.46 1.20 5.73 1.98
Fe 1.58 1.09 1.64 1.29 6.92 1.42
Co 1.77 1.31 1.76 1.52 7.94 0.95
Ni 1.80 1.82 1.89 1.89 9.03 0.48
Supplementary Note 8 Spin fluctuation in Kondo impurities
The spin fluctuation in all magnetic impurities can be better understood by calculating the spin-spin correlation within the $3d$ shell
$S_{ij}=\langle\hat{S}_{i}\hat{S}_{j}\rangle-\langle\hat{S}_{i}\rangle\langle\hat{S}_{j}\rangle.$ (S11)
As shown in Fig. S6, Cr has the strongest intra- and inter-orbital spin correlation among all impurities. From Cr to Co, the intra-orbital spin correlation in $t_{2g}$ orbitals is largely suppressed, while the intra- and inter-orbital spin correlation within the $e_{g}$ orbitals stays strong or moderate, which is consistent with the orbital-dependent trend in quasiparticle renormalization. Fig. S6 also reveals that, in the Mn, Fe, and Co systems, the inter-orbital spin correlations have the following relation: $S_{i\neq j}(t_{2g}\text{-}t_{2g})<S_{i\neq j}(t_{2g}\text{-}e_{g})<S_{i\neq j}(e_{g}\text{-}e_{g})$.
Figure S6: Orbital-resolved spin-spin correlations for Cr, Mn, Fe, and Co impurities.
Supplementary Note 9 Effective $3d$-model calculations
To compare our all-orbital DMFT simulations against downfolded effective model calculations, we derived a multi-orbital Anderson impurity model with Hamiltonian
$H=\sum_{ij\sigma}\tilde{F}_{ij}f^{\dagger}_{i\sigma}f_{j\sigma}+\frac{1}{2}\sum_{ijkl}\sum_{\sigma\sigma^{\prime}}U_{ijkl}f^{\dagger}_{i\sigma}f^{\dagger}_{j\sigma^{\prime}}f_{l\sigma^{\prime}}f_{k\sigma}+\sum_{k\sigma}\epsilon_{k}c^{\dagger}_{k\sigma}c_{k\sigma}+\sum_{ki\sigma}V_{ik}(f^{\dagger}_{i\sigma}c_{k\sigma}+c^{\dagger}_{k\sigma}f_{i\sigma}),$ (S12)
where the indices $i,j,k,l$ run over five $3d$ orbitals and $U_{ijkl}$ is the screened Coulomb interaction tensor within the $3d$ shell. In practice, approximations such as the density-density approximation are often employed for the Coulomb interaction tensor. Here, we adopted a Kanamori Hamiltonian [Valli2020] that goes beyond the density-density approximation, where the Coulomb tensor (second term in Eq. S12) is written as
$H_{K}=\sum_{i}U_{iiii}\hat{n}_{i\uparrow}\hat{n}_{i\downarrow}+\sum_{i\neq j}\sum_{\sigma\sigma^{\prime}}(U_{ijij}-U_{ijji}\delta_{\sigma\sigma^{\prime}})\hat{n}_{i\sigma}\hat{n}_{j\sigma^{\prime}}+\sum_{i\neq j}U_{ijji}(f^{\dagger}_{i\uparrow}f^{\dagger}_{j\downarrow}f_{i\downarrow}f_{j\uparrow}-f^{\dagger}_{i\uparrow}f^{\dagger}_{i\downarrow}f_{j\uparrow}f_{j\downarrow}).$ (S13)
We give the Coulomb integrals for $3d$ electrons in the basis of cubic harmonics, where $U_{iiii}$, $U_{ijij}$, and $U_{ijji}$ are expressed using parameters $U_{0},J_{1},J_{2},J_{3},J_{4}$. The readers are referred to Ref. [Valli2020] for the detailed parametrization. The $U_{0},J_{1},J_{2},J_{3},J_{4}$ parameters can be expressed in terms of Slater integrals $F^{0}$, $F^{2}$, and $F^{4}$:
$U_{0}$ $=F^{0}+\frac{8}{7}\frac{1}{14}(F^{2}+F^{4}),$ (S14)
$J_{1}$ $=\frac{1}{49}(3F^{2}+\frac{20}{9}F^{4}),$ (S15)
[…]
The Slater integrals $F^{0}$, $F^{2}$, and $F^{4}$ are obtained from two parameters $U$ and $J$, where $U=F_{0}$ and $J=\frac{1}{14}(F^{2}+F^{4})$ with a constant ratio $F^{2}/F^{4}=0.625$. In summary, the Kanamori Hamiltonian in Eq. S13 is fully characterized by two parameters $U$ and $J$. In Kondo simulation literature, $U$ and $J$ are usually treated as adjustable parameters. Here, we took the $U$ and $J$ values of magnetic impurities from cRPA calculations in Ref. [Sasioglu2011], which are listed in Table S6.
Table S6: Coulomb interaction parameters in Kanamori Hamiltonian taken from Ref. [Sasioglu2011].
Parameter (eV) Ti V Cr Mn Fe Co Ni
$U$ 3.1 3.2 4.4 4.4 3.8 4.3 3.8
$J$ 0.5 0.6 0.7 0.7 0.7 0.8 0.8
Table S7: Quasiparticle renormalization weights $Z$ and Kondo temperatures $T_{\mathrm{K}}$ of magnetic impurities in bulk Cu from five-orbital model Hamiltonian calculations.
Impurity $Z(t_{2g})$ $Z(e_{g})$ $T_{\mathrm{K}}(t_{2g})$ (K) $T_{\mathrm{K}}(e_{g})$ (K)
Ti 0.92 0.95 4953 5556
V 0.63 0.87 2993 4431
Cr 0.31 0.55 1318 2445
Mn 0.11 0.062 449 243
Fe 0.45 0.11 1769 398
Co 0.71 0.72 2800 2524
Ni 0.99 0.99 3999 3383
Table S8: Orbital occupancies $n$ and natural orbital occupancies $n_{\mathrm{nat}}$ of magnetic impurities in bulk Cu from five-orbital model Hamiltonian calculations.
Impurity $n(t_{2g})$ $n(e_{g})$ $n_{\mathrm{nat}}(t_{2g})$ $n_{\mathrm{nat}}(e_{g})$ $n(\mathrm{3d})$
Ti 0.32 0.13 1.98 1.94 1.23
V 0.75 0.15 1.86 1.89 2.54
Cr 1.23 0.27 1.63 1.85 4.23
Mn 1.26 0.84 1.59 1.50 5.45
Fe 1.60 1.07 1.84 1.52 6.95
Co 1.68 1.74 1.93 1.80 8.51
Ni 1.94 1.97 1.91 1.88 9.76
The one-particle Hamiltonian in Eq. S12 was treated in a similar fashion as in the all-orbital DMFT calculations, where we adopted the Hartree-Fock effective Hamiltonian for five $3d$ orbitals so that the double-counting term can be exactly removed. The $3d$ block of the DFT hybridization function was discretized on the same non-uniform grid to obtain the bath parameters $\epsilon_{k}$ and $V_{ik}$, which resulted in 49 bath orbitals per $3d$ impurity orbital.
To solve the embedding problem consisting of 5 impurity orbitals and 245 bath orbitals, we employed the same active-space DMRG solver. A Hartree-Fock calculation with fixed chemical potential at the DFT level was first performed, followed by a CISD calculation on the full embedding problem. A (40e, 40o) natural-orbital active space was derived by diagonalizing the CISD density matrix. We then conducted ground-state DMRG calculations with bond dimension $M=3500$ on the (40e, 40o) active space and further derived a (20e, 20o) DMRG natural-orbital active space. Finally, a dynamical DMRG calculation was done on the (20e, 20o) active space with bond dimension $M=1200$, and DMRG self-energies were used to estimate the quasiparticle renormalizations and Kondo temperatures in magnetic impurities. We summarize this result in Table S7 and the ground-state properties in Table S8. We find that the natural occupancies in downfolded model calculations are less fractional than the results in all-orbital DMFT calculations. In the meantime, the quasiparticle renormalization weights and the predicted Kondo temperatures are much higher in the model calculations compared to all-orbital DMFT.